OpenStudy (shadow-the-hedgehog):

A circular piece of fabric has a radius of 1.2 feet. The fabric sells for $5.40 per square foot. What is the total cost of the circular piece of fabric? A. $6.48 B. $20.35 C. $24.42 D. $40.69

OpenStudy (ltrout):

First find the area of the circle. A = pi * r^2. A = pi * 1.2^2 A = 4.52389342117ft^2 Now multiply that by the price of 5.40 / sqft. 4.52389342117 * 5.40 = 24.42$

OpenStudy (shadow-the-hedgehog):

Thanks